OpenStudy (anonymous):

The product of 3 and a number is 12 more than the sum of the number and 26. What is the number?

OpenStudy (anonymous):

How would you write this as an equation?

OpenStudy (anonymous):

2 x X =12 X =?

OpenStudy (anonymous):

@nelsonjedi

OpenStudy (anonymous):

@One098

OpenStudy (anonymous):

@Here_to_Help15 @chosenmatt

OpenStudy (anonymous):

3x-12=x+26

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

Somenone please help

OpenStudy (anonymous):

@Marki

OpenStudy (anonymous):

so u would have 2x=38

OpenStudy (anonymous):

x=38/2

OpenStudy (anonymous):

19

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

2 * 19?

OpenStudy (anonymous):

So the answer is 19? Is that what I put? :)

OpenStudy (anonymous):

haha yeah lol 19 is correct